OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

@pooja195 nope, what are the steps I can probably figure it out

OpenStudy (calculusxy):

you can use one of two methods, elimination or substitution

OpenStudy (calculusxy):

you would want to find the value of y because it will help you, once you substitute that y value back in to the second equation, to find x

OpenStudy (calculusxy):

\(x - y = -3\) \(-y = -3 - x\) \( y = 3 + x\) substitute \(y = 3 + x \) to where y is in the second equation \(x + 3(3 + x) = 5\) \(x + 9 + 3x = 5\) \(4x + 9 = 5\) \(4x + 9 - 9 = 5 - 9\) \(4x = -4\) \(x = -1\)

OpenStudy (calculusxy):

did you understand?

OpenStudy (wolf1728):

Here's another way to solve it: x − y = −3 x + 3y = 5 multiply the bottom equation by -1 x -y = -3 -x -3y = -5 add both equations -4y = -8 y = 2

OpenStudy (wolf1728):

redsky - do you understand either method?